DX AO Observations

07:50
  • Log Started
08:21
  • Preset
  • 100 Hz
  • Bin 4
  • Mag 16
  • 36 Modes
  • Datasets:
  • Gain 3.4 1.94 1.90
  • Dataset: 20170330_082610 Probably incomplete.
08:39
  • Adsec/stop/start performed because processes disconnected, also restarted w_unit for same reason. Moxa was still on after restart?!
08:44
  • New Preset
  • Bin 4
  • 100 Hz
  • 36 Modes
  • Gain: 3.40 2.0 1.76
  • Dataset: 20170330_085252 Odd behavior in the modal plot
  • Dataset: 20170330_085844 2000 frames
09:42
  • SX failed to resume after offset.
  • DX was successful.
  • Closed loop we get magnitude 15.7. The stated magnitude is 14.7, so we are about 1 magnitude off.
  • Dataset: 20170330_100003 2000 frames
  • Dataset: 20170330_100046 4000 frames
10:03
  • New Target
  • Bin 2
  • 153 Modes
  • 625 Hz
  • Mag 12.1
  • Gain 2.55 1.1 0.62
  • Dataset: 20170330_101208
10:47
  • SX failed again.
  • Restarted same preset
  • Bin 2
  • 153 Modes
  • 625 Hz
  • Mag 12.2
  • Gain 1.7 1.26 0.60
11:25
  • Offset Failed on DX this time.
  • Reclosed loop with gains 1.7 1.25 0.60
12:17
  • NCPA check
  • No NCPA
  • K-Short Fiter
  • Luci Image 101-110
  • Dataset: 20170330_121805
  • Optical gain 0.12
  • NCPA ON
  • Luci Images 111-120
  • Dataset: 20170330_122045
12:27
  • Bright Star NCPA test
  • No NCPA
  • Bin 1 990 Hz
  • 300 Modes
  • Gains 1.7 1.2 0.6
  • Dataset: 20170330_123600
  • Luci Images: 121
  • NCPA On
  • Optical Gain 0.17
  • Dataset: 20170330_123935
  • Luci Image 122
  • Dataset: 20170330_124531
  • Luci Image 123, 124 FeII 125
  • NCPA Off
  • Dataset: 20170330_124808
  • Luci Image 126

-- %USERSIG{GregoryTaylor - 2017-03-30}%
Topic revision: r2 - 30 Mar 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback